# Simultaneous Masking, Not Prompting Optimization: A Paradigm Shift in Fine-tuning LLMs for Simultaneous Translation

### v1

@@ Abstract
Large language models (LLMs) have achieved state-of-the-art performance in various language processing tasks, motivating their adoption in simultaneous translation. Current fine-tuning methods to adapt LLMs for simultaneous translation focus on prompting optimization strategies using either data augmentation or prompt structure modifications. However, these methods suffer from several issues, such as an unnecessarily expanded training set, computational inefficiency from dumping the KV cache, increased prompt sizes, or restriction to a single decision policy. To eliminate these issues, we propose a new paradigm in fine-tuning LLMs for simultaneous translation, called SimulMask. It utilizes a novel attention mask technique that models simultaneous translation during fine-tuning by masking attention connections under a desired decision policy. Applying the proposed SimulMask on a Falcon LLM for the IWSLT 2017 dataset, we have observed a significant translation quality improvement compared to state-of-the-art prompting optimization strategies on three language pairs when averaged across four different latency regimes while reducing the computational cost.
@@ 1 Introduction
Simultaneous translation refers to the process of producing a target output translation concurrently with an oncoming source input. In our increasingly interconnected world, where communication across languages in real-time is desired, simultaneous translation is becoming a requirement. Unfortunately, the task of simultaneous translation is extremely straining for human interpreters, and among the limited number of simultaneous human interpreters, their maximum translation length is around thirty minutes Moser-Mercer et al. (1998). As such, there is a need for machine learning models to alleviate the burden placed on simultaneous human interpreters.
@@
Current literature has primarily focused on adapting end-to-end Transformer models Vaswani et al. (2017) to overcome the difficulties of simultaneous translation due to their reduced parameter counts and greater inference speedMa et al. (2020b). However, the recent successes of large language models (LLMs) Touvron et al. (2023); Jiang et al. (2023); Almazrouei et al. (2023) has prompted preliminary research applying them to simultaneous translation through fine-tuning and inference techniques Agostinelli et al. (2023); Wang et al. (2023); Koshkin et al. (2024); Wang et al. (2024); Guo et al. (2024). Unfortunately, most modern works have neglected the computational increases created by dumping the target sequence’s key and value (KV) cache Wang et al. (2024). Furthermore, there has yet to be a universal approach to fine-tuning LLMs for simultaneous translation that is not unnecessarily computationally expensive by either expanding the dataset through data augmentation, a process referred to as prefix fine-tuning Agostinelli et al. (2023); Wang et al. (2023); Koshkin et al. (2024) or increasing the prompt length through prompt restructuring Koshkin et al. (2024); Wang et al. (2024).
@@
The lack of an efficient fine-tuning strategy of LLMs for simultaneous translation has led us to propose a new paradigm, referred to as SimulMask. SimulMask is a novel attention mask to model simultaneous translation during fine-tuning by redistributing the attention connections under a decision policy. By design, SimulMask is broadly applicable to both flexible and fixed decision policies, creating a path forward for future work to build upon it. Furthermore, if we eliminate injecting positional information into the keys and values through a modified ALiBi Press et al. (2021), SimulMask allows for key and value caching during simultaneous translation without accuracy degradation.
To validate the efficacy of SimulMask, we fine-tuned and evaluated 1.3 billion parameter Falcon models pre-trained on the RefinedWeb dataset Almazrouei et al. (2023); Penedo et al. (2023) using the IWSLT 2017 dataset Cettolo et al. (2017). All models were fine-tuned using a wait-k decision policy. From the averaged evaluation results across all wait-k values, we found an LLM fine-tuned with SimulMask and evaluated with KV caching outperforms an LLM fine-tuned with prefix fine-tuning and evaluated without KV caching by 7.34, 3.00, and 2.98 BLEU on the English-French, English-Dutch, and English-Italian language pairs, respectively. Furthermore, on average, with the wait-3 policy, recomputing the KV cache contributes toward 26.8 % of the computation time and 87.9 % of the FLOPs to generate a target translation on an A40 GPU, substantiating the need for SimulMask, which enables inference time KV caching to avoid those computations.
In summary, the main contributions of the paper include:
@@
    Providing insights on the shortcomings of current literature in adapting LLMs to simultaneous translation.
@@
    Proposing a novel attention masking approach to fine-tune LLMs for simultaneous translation that allows for efficient training and inference.
@@
    Demonstrating the efficacy of our proposed approach in terms of translation quality and computational costs by evaluating them on multiple language pairs across many wait-k values.
@@ 2.1 Masked Transformer Self-Attention
We briefly review self-attention functionality in transformers (Vaswani et al., 2017) focusing on masking behavior in Equation 1, which describes per-head attention calculations for some attention head $h$. For some input ${\bm{X}}\in\mathbb{R}^{N\times d_{model}}$ with $N$ being the sequence length and $d_{model}$ representing model dimensionality, the query matrix ${\bm{Q}}^{(h)}={\bm{X}}{\bm{W}}^{(h)}_{q}$, the key matrix ${\bm{K}}^{(h)}={\bm{X}}{\bm{W}}^{(h)}_{k}$, and the value matrix ${\bm{V}}^{(h)}={\bm{X}}{\bm{W}}^{(h)}_{v}$. The aforementioned weight matrices for some attention head $h$ are defined as ${\bm{W}}^{(h)}_{q}\in\mathbb{R}^{d_{model}\times d_{head}}$, ${\bm{W}}^{(h)}_{k}\in\mathbb{R}^{d_{model}\times d_{head}}$ and ${\bm{W}}^{(h)}_{v}\in\mathbb{R}^{d_{model}\times d_{head}}$, with $d_{head}$ representing per-attention head dimensionality. We define each individual query, key, and value in their respective matrices as ${\bm{q}}_{i}^{(h)}\in{\bm{Q}}^{(h)}$, ${\bm{k}}_{i}^{(h)}\in{\bm{K}}^{(h)}$, and ${\bm{v}}_{i}^{(h)}\in{\bm{V}}^{(h)}$. ${\bm{M}}$ is defined as an optional attention mask.
${\bm{A}}^{(h)}=\texttt{softmax}\left(\frac{{\bm{Q}}^{(h)}{{\bm{K}}^{(h)}}^{T}+{\bm{M}}}{\sqrt{d_{head}}}\right){\bm{V}}^{(h)}$ (1)
Critically, ${\bm{M}}$ functions by modeling context limitations or time-based dependencies that might exist during inference but do not exist during training/fine-tuning. For generative transformer blocks or architectures, ${\bm{M}}$ is defined as a causal attention mask, where each entry, ${\bm{M}}_{ij}$ is represented by Equation 2.
@@ 2.2 Simultaneous Translation
Simultaneous translation is dictated by read-write decision policies, whereby a model will wait a specific amount of time before alternating between reading and writing in fixed or flexible intervals. One fixed decision policy for simultaneous translation that is broadly adopted as a common baseline to build on due to its effectiveness and simplicity is the wait-k policy Ma et al. (2019). As the name suggests, the wait-k policy will wait for k words before alternating between writing and reading a word. Although effective in simultaneous translation, alternative adaptive policies have gained traction, which base reading and writing on an auxiliary model or a predefined set of rules Cho and Esipova (2016); Gu et al. (2017); Zheng et al. (2019). While capable of impressive results, such adaptive policies often incur additional computational costs and delays.
A Transformer is trained for a simultaneous translation decision policy by masking attention scores in the encoder self-attention and the decoder cross-attention. In the case of the encoder self-attention, each source token is prevented from attending to future source tokens following a decision policy Ma et al. (2019). For example, if the source sequence length was 5 tokens and the first read step reads 2 tokens, the second read step reads 1 token, and the third read step reads 2 tokens, then the respective attention mask is provided in Figure 1.
@@
In Equation 3, $f(t)$, is a decision policy function that denotes the cumulative number of source hidden states to read when predicting target token $t$. Each source hidden state aligns with one source input token.
2.3 Applying Large Language Models for Simultaneous Translation
LLMs have demonstrated remarkable performance on the task of neural machine translation Moslem et al. (2023); Vilar et al. (2023); Xu et al. (2023); Zhang et al. (2023); Iyer et al. (2023). Such successes have prompted recent works to extend the reach of LLMs into the realm of simultaneous translation Agostinelli et al. (2023); Wang et al. (2023); Koshkin et al. (2024); Wang et al. (2024); Guo et al. (2024). LLMs are especially promising for the field of simultaneous translation due to their strong understanding of language semantics and meaning. Intuitively, SimulMT LLMs inject holistic linguistic knowledge that could allow for correct translation decisions when facing difficult contextual obstacles (e.g., translating a verb in a target language without access to that verb in the source language).
Unfortunately, Equation 3 is no longer effective in modeling simultaneous translation for decoder-only LLMs as with the decoder of a Transformer. The reason is Equation 3 is constructed specifically for the cross-attention calculation between keys exclusively from the source and queries exclusively from the target, as in Transformers. Since LLMs perform a self-attention, including the prompt, the source, and the target, Equation 3 cannot properly mask the source from the target with the additional prompt and target sequences included in the keys, and the additional prompt and source sequences included in the queries. Furthermore, Equation 3 does not enforce the autoregressive language modeling behavior of LLMs. As such, alternative means to model simultaneous translation have been proposed, leveraging prompting optimization.
@@ 3 Prompting Optimization Methods
Current methods of fine-tuning LLMs for simultaneous translation fall into the general category of prompting optimization. We define prompting optimization as either employing data augmentation to help with prompting Koshkin et al. (2024); Wang et al. (2023); Agostinelli et al. (2023) or redefining the prompt structure Wang et al. (2024); Koshkin et al. (2024) to somewhat simulate simultaneous translation. In this section, we will cover these prompting optimization strategies in depth.
@@ 3.1 Data Augmentation
Prompting optimization focusing on data augmentation resorts to subdividing each sentence in a dataset into multiple partial sentence pairs. These partial sentence pairs mimic simultaneous translation, as simultaneous translation produces outputs with a partial input. We label such a method as prefix fine-tuning, and although the high-level procedure is identical amongst current works, the algorithms employed to obtain these partial sentence pairs are unique. In the case of Agostinelli et al. (2023), each source-target sentence pair is subdivided according to the wait-k policy such that if we order the new samples from smallest to largest, each subsequent sentence pair will have one additional target word and source word so long as the end of the target or source is not reached. Upon completion there will be $\texttt{max}(|S|-(k-1),|T|)$ sentence pairs, where $|S|$ and $|T|$ are the original source and target sequence lengths. The approach requires the model to predict only the final target word in the sequence during fine-tuning.
Alternatively, Wang et al. (2023) randomly sampled a subset of sentence pairs from the dataset and truncated the source sentence to be 20% to 80% of the full length according to a uniform distribution. They obtained the respective target translations by prompting ChatGPT. The new truncated source-target sentence pairs were then added to the complete dataset to expand the overall fine-tuning dataset.
@@ 4 Analysis and Shortcomings of Prompting Optimization Methods
Prompting optimization, while works to a certain degree, is inherently deficient, possessing a host of fine-tuning and inference issues. These issues include a persistent fine-tuning-inference mismatch, consistent positional confusion in the target sequence, and high computational costs. We will first present the fine-tuning-inference mismatch, followed by the positional confusion problem, and finish with the computational burdens facing the current paradigm.
@@ 4.1 Fine-tuning/Inference Mismatch
A fine-tuning-inference mismatch is a mismatch between an LLM’s fine-tuning and inference environments. For instance, fine-tuning an LLM for neural machine translation where the entire sentence is available and deploying it for simultaneous translation where little of the sentence is available when beginning generation will create a massive inference time fine-tuning-inference mismatch. Furthermore, the LLM must be fine-tuned to accommodate KV caching, the process of caching the keys and values at inference to prevent recomputation. Overall, fine-tuning for simultaneous translation aims to minimize the fine-tuning-inference mismatch.
Unfortunately, prefix fine-tuning precludes high-quality simultaneous translation with KV caching Agostinelli et al. (2023); Koshkin et al. (2024); Wang et al. (2023) as with the continuously increasing prompt size, each subsequent entry in the KV cache continuously becomes outdated, a scenario the LLM was not fine-tuned to handle. For example, suppose we have an LLM in the middle of simultaneous translation using KV caching adhering to a wait-1 policy with the following prompting structure: “Translate the following sentence from English to German: $s_{1},s_{2},...,s_{i}$ [a]: $t_{1},t_{2},...,t_{i}$”. Now suppose we read $s_{i+1}$. Then, in the subsequent write decision, the LLM must produce $t_{i+1}$. At the given write step, the KV cache entries for $t_{1},t_{2},...,t_{i}$ are each generated with a different subset of the source sequence $s_{1},s_{2},...,s_{i}$. For instance, the keys and values that generated $t_{1}$ could only attend to $s_{1}$, whereas the keys and values of $t_{i}$ could attend to $s_{1},s_{2},...,s_{i}$. However, during prefix fine-tuning, the LLM is fine-tuned to predict $t_{i+1}$ as if the KV cache for $t_{1},t_{2},...,t_{i}$ are each generated with the same subset of the source sequence $s_{1},s_{2},...,s_{i}$. As such, the KV cache for $t_{1}$ and $t_{i}$ would both be generated by attending to the source sequence $s_{1},s_{2},...,s_{i}$. Such fine-tuning-inference mismatch is unsolved through conventional prompting structures and is necessitated to increase translation quality.
Prompting restructuring also creates additional fine-tuning-inference mismatches. In Koshkin et al. (2024); Wang et al. (2024), they all fine-tune for an oracle decision policy. However, at inference, such an oracle decision policy is not truly achievable, creating a mismatch. Furthermore, since the LLMs that leverage prompt restructuring encapsulate a specific oracle decision policy into their fine-tuning curriculum, extending them to alternative decision policies at inference is infeasible without incurring a mismatch. As such, there is a need for a flexible method adaptable to a range of decision policies that also eliminates the fine-tuning-inference mismatch.
@@ 4.2 Positional Confusion
Positional confusion describes the process whereby the relative and/or global positional information during simultaneous translation progressively becomes incorrect. Unfortunately, most simultaneous translation LLMs using KV caching suffer from this positional confusion Agostinelli et al. (2023); Koshkin et al. (2024); Wang et al. (2023). The reason is that as the source sequence grows with simultaneous translation, the target sequence shifts, necessitating the target sequence’s positional information to follow suit. However, since KV caching is employed, updating the keys and values is infeasible, causing them to hold incorrect positional information.
Aligning with our previous example, the sequence portion that would experience this positional confusion would be “[a]: $t_{1},t_{2},...,t_{i}$”. For example, after the first prediction step, the positional distance between $s_{1}$ and [a]: and $t_{1}$ would be 1 and 2, respectively. Then, after the second read, where the source sequence is $s_{1},s_{2}$ the positional distance between $s_{1}$ and [a]: and $t_{1}$ would change to 2 and 3, respectively. However, while using KV caching, this positional distance would remain 1 and 2 in the keys and/or values for subsequent predictions, causing positional confusion. Continuing translation would see the gap between the true positional distance and the positional distance in the KV cache grow. Identifying an effective method to deal with positional confusion is a requirement to prevent LLM hallucinations.

### v4

@@ Abstract
Large language models (LLMs) have achieved state-of-the-art performance in various language processing tasks, motivating their adoption in simultaneous translation. Current fine-tuning methods to adapt LLMs for simultaneous translation focus on prompting optimization strategies using either data augmentation or prompt structure modifications. However, these methods suffer from several issues, such as unnecessarily expanded training sets, computational inefficiency from dumping the key and value cache, increased prompt sizes, or restriction to a single decision policy. To eliminate these issues, in this work, we propose SimulMask, a new paradigm for fine-tuning LLMs for simultaneous translation. It utilizes a novel attention mask approach that models simultaneous translation during fine-tuning by masking attention for a desired decision policy. Applying the proposed SimulMask on a Falcon LLM for the IWSLT 2017 dataset, we have observed a significant translation quality improvement compared to state-of-the-art prompting optimization strategies on five language pairs while reducing the computational cost.
@@ 1 Introduction
Simultaneous translation refers to the process of producing a target output translation concurrently with an oncoming source input. In our increasingly interconnected world, where communication across languages in real-time is desired, simultaneous translation is becoming a requirement. As such, there is a need for machine learning models to fill this role.
@@
Current literature has primarily focused on adapting end-to-end Transformer models Vaswani et al. (2017) to overcome the difficulties of simultaneous machine translation (SimulMT) due to their reduced parameter counts and greater inference speedMa et al. (2020b). However, the recent successes of large language models (LLMs) Touvron et al. (2023); Jiang et al. (2023); Almazrouei et al. (2023) has prompted preliminary research applying them to SimulMT through fine-tuning and inference techniques Agostinelli et al. (2023); Wang et al. (2023); Koshkin et al. (2024); Wang et al. (2024); Guo et al. (2024). Unfortunately, most modern works have neglected the computational increases created by dumping the target sequence’s key and value (KV) cache Wang et al. (2024). Furthermore, there has yet to be a universal approach to fine-tuning LLMs for SimulMT that is not unnecessarily computationally expensive by either expanding the dataset through data augmentation, a process referred to as prefix fine-tuning Agostinelli et al. (2023); Wang et al. (2023); Koshkin et al. (2024) or increasing the prompt length through prompt restructuring Koshkin et al. (2024); Wang et al. (2024).
@@
The lack of an efficient fine-tuning strategy of LLMs for SimulMT has led us to propose a new paradigm, referred to as SimulMask. SimulMask is a novel attention mask to model SimulMT during fine-tuning by redistributing the attention under a decision policy. By design, SimulMask is broadly applicable to both flexible and fixed decision policies, creating a path forward for future work to build upon it. Furthermore, if we avoid injecting positional information into the keys and values through a modified ALiBi Press et al. (2021), SimulMask allows for KV caching during SimulMT without accuracy degradation.
To validate the efficacy of SimulMask, we fine-tuned and evaluated 1.3 billion parameter Falcon models pre-trained on the RefinedWeb dataset using SimulMask Almazrouei et al. (2023); Penedo et al. (2023) and compared them against identical Falcon models that adopt existing prefix fine-tuning or prompt restructuring methods on the IWSLT 2017 dataset Cettolo et al. (2017). From the results, we demonstrate models fine-tuned with SimulMask outperform prefix fine-tuning and prompt restructuring models at SimulMT for a given latency regime with a reduced computational cost.
The main contributions of the paper include:
@@
    Providing insights on the shortcomings of current methods in adapting LLMs to SimulMT.
@@
    Proposing a novel attention masking approach to fine-tune SimulMT LLMs that enables efficient training and inference.
@@
    Demonstrating the efficacy of our approach in terms of translation quality and computational costs by evaluating them on multiple language pairs across varied latencies.
@@ 2.1 Masked Transformer Self-Attention
We briefly review self-attention functionality in Transformers Vaswani et al. (2017) focusing on masking behavior in Equation 1. ${\bm{M}}$ is defined as an optional attention mask.
${\bm{A}}=\text{softmax}\left(\frac{{\bm{Q}}{\bm{K}}^{T}+{\bm{M}}}{\sqrt{d_{head}}}\right){\bm{V}}$ (1)
Critically, ${\bm{M}}$ functions by modeling context limitations or time-based dependencies that might exist during inference but do not exist during training/fine-tuning. For generative Transformer blocks or architectures, ${\bm{M}}$ is defined as a causal attention mask, where each entry, ${\bm{M}}_{ij}$, is represented by Equation 2 to avoid attending to the future.
@@ 2.2 Simultaneous Translation
SimulMT is dictated by read-write decision policies, whereby a model will wait a specific amount of time before alternating between reading and writing in fixed or flexible intervals. One fixed decision policy for SimulMT that is broadly adopted as a common baseline to build on due to its effectiveness and simplicity is the wait-k policy Ma et al. (2019). As the name suggests, the wait-k policy will wait for k words before alternating between writing and reading a word. Although effective in SimulMT, alternative adaptive policies have gained traction, which base reading and writing on an auxiliary model or a predefined set of rules Cho and Esipova (2016); Gu et al. (2017); Zheng et al. (2019). While capable of impressive results, such adaptive policies often incur additional computational costs.
A Transformer is trained for a SimulMT decision policy by masking attention scores in the encoder self-attention and the decoder cross-attention. In the case of the encoder self-attention, each source token is prevented from attending to future source tokens following a decision policy Ma et al. (2019). An example mask is provided in Appendix A.1.
@@
In Equation 3, $f(t)$, is a decision policy function that denotes the cumulative number of source hidden states to read when predicting target token $t$.
2.3 Applying LLMs to SimulMT
LLMs have demonstrated remarkable performance on neural machine translation (NMT) Moslem et al. (2023); Vilar et al. (2023); Xu et al. (2023); Zhang et al. (2023); Iyer et al. (2023). Such successes have prompted recent works to extend the reach of LLMs into the realm of SimulMT Agostinelli et al. (2023); Wang et al. (2023); Koshkin et al. (2024); Wang et al. (2024); Guo et al. (2024). LLMs are especially promising for the field of SimulMT due to their strong understanding of language semantics and meaning. Intuitively, SimulMT LLMs inject holistic linguistic knowledge that could allow for correct translation decisions when facing difficult contextual obstacles (e.g., translating a verb in a target language without access to that verb in the source language).
Unfortunately, Equation 3 is no longer effective in modeling SimulMT for decoder-only LLMs as with the decoder of a classical Transformer. The reason is that Equation 3 is constructed specifically for the cross-attention calculation between keys exclusively from the source and queries exclusively from the target, as in Transformers. In contrast, LLMs perform self-attentions involving the prompt, the source, and the target concurrently. Equation 3 can no longer properly mask the source (keys) from the target (queries), due to the additional prompt and target sequences in the keys and the additional prompt and source sequences in the queries. Furthermore, Equation 3 does not enforce the autoregressive language modeling behavior of LLMs. As such, alternative means to model SimulMT have been proposed, leveraging prompting optimization.
@@ 3 Prompting Optimization Methods
Current methods of fine-tuning LLMs for SimulMT fall under prompting optimization. We define prompting optimization as either employing data augmentation to help with prompting Koshkin et al. (2024); Wang et al. (2023); Agostinelli et al. (2023) or redefining the prompt structure Wang et al. (2024); Koshkin et al. (2024) to somewhat simulate SimulMT.
@@ 3.1 Data Augmentation
Prompting optimization focusing on data augmentation resorts to subdividing each sentence in a dataset into multiple partial sentence pairs. These partial sentence pairs mimic SimulMT, as SimulMT produces outputs with a partial input. We label such a method as prefix fine-tuning, and although the high-level procedure is identical amongst current works, the algorithms employed to obtain these partial sentence pairs are unique. In the case of Agostinelli et al. (2023), each source-target sentence pair is subdivided according to the wait-k policy such that if we order the new samples from smallest to largest, each subsequent sentence pair will have one additional target word and source word so long as the end of the target or source is not reached. Upon completion there will be $\texttt{max}(|S|-(k-1),|T|)$ sentence pairs, where $|S|$ and $|T|$ are the original source and target sequence lengths. The approach requires the model to predict only the final target word in the sequence during fine-tuning.
Alternatively, Wang et al. (2023) randomly sampled a subset of sentence pairs from the dataset and truncated the source sentence to be 20% to 80% of the full length according to a uniform distribution. They obtained the respective target translations by prompting ChatGPT (gpt-3.5-turbo). The new truncated source-target sentence pairs were then added to the complete dataset, expanding it.
@@ 4 Analysis and Shortcomings of Prompting Optimization Methods
Prompting optimization, while functional to a certain degree, is inherently deficient, possessing a host of fine-tuning and inference issues. These issues include a persistent fine-tuning-inference mismatch, consistent positional confusion in the target sequence, and high computational costs.
@@ 4.1 Fine-tuning/Inference Mismatch
A fine-tuning-inference mismatch is a mismatch between a LLM’s fine-tuning and inference environments. For instance, fine-tuning a LLM for NMT where the entire sentence is available and deploying it for SimulMT where little of the sentence is available when beginning generation will create a massive inference time fine-tuning-inference mismatch. Furthermore, the LLM must be fine-tuned to accommodate KV caching, the process of caching the keys and values at inference to prevent recomputation. Overall, fine-tuning for SimulMT aims to minimize the fine-tuning-inference mismatch.
[…]
Prompting restructuring also creates additional fine-tuning-inference mismatches. In Koshkin et al. (2024) and Wang et al. (2024), they all fine-tune for an oracle decision policy. However, at inference, such an oracle decision policy is not truly achievable, creating a mismatch. Furthermore, since the LLMs that leverage prompt restructuring encapsulate a specific oracle decision policy into their fine-tuning curriculum, extending them to alternative decision policies at inference is infeasible without incurring a mismatch. This calls for a new, flexible method adaptable to a range of decision policies while also eliminating the fine-tuning-inference mismatch.
@@ 4.2 Positional Confusion
Positional confusion describes the process whereby the relative and/or global positional information during SimulMT progressively becomes incorrect. Unfortunately, most SimulMT LLMs using KV caching suffer from this positional confusion Agostinelli et al. (2023); Koshkin et al. (2024); Wang et al. (2023). The reason is that as the source sequence grows with SimulMT, the target sequence also shifts, necessitating the target sequence’s positional information to follow suit. However, since KV caching is employed, the positional information held in the keys and values is not properly updated.
Aligning with our previous example, for the sequence portion “[a]: $t_{1},t_{2},...,t_{i}$”, after the first prediction step, the positional distance between $s_{1}$ and [a]: and between $s_{1}$ and $t_{1}$ would be 1 and 2, respectively. Then, after the second read, where the source sequence is $s_{1},s_{2}$, the positional distance between $s_{1}$ and [a]: and $t_{1}$ would change to 2 and 3, respectively. However, while using KV caching, this positional distance would remain 1 and 2 in the keys and/or values for subsequent predictions, causing positional confusion. Continuing translation would see the increased gap between the true positional distance and the stale positional distance in the KV cache. Therefore, we need to identify an effective method to deal with positional confusion that is essential to prevent LLM hallucinations.